### v1

@@ Abstract
We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of spatial autoregressive models. Our theory covers error dependence of a general form, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial coefficients. One application of our method finds evidence for constant returns to scale in the production function of the Chinese nonmetal mineral industry, while another finds a nonlinear impact of the distance to the employment center on housing prices in Boston. A simulation study confirms that our tests perform well in finite-samples.
@@
Keywords: Spatial autoregression, varying coefficients, inference
@@ 1.   Introduction
This paper develops a general framework for inference on varying coefficient spatial autoregressive (SAR) models that is as easy to implement in practice as familiar Wald tests. SAR models are fast becoming a workhorse for researchers working in settings where economic agents interact via economic or social links. Such links can be geographic proximity, or proximity in some more general sense such as a social network. The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal. No surprise then that the baseline SAR model, due to Cliff and Ord (1981), has become much studied by econometricians. This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998) and Lee (2004).
@@
In keeping with the growing importance of SAR models, various strands of the literature have explored how to widen their scope. One such is the focus on varying coefficient SAR models, wherein model parameters are allowed to be unknown functions of some observed economic variable(s). This permits a flexible effect of the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g. Fan and Zhang, 1999, 2008). In the SAR context, varying coefficient models have been studied by Sun et al. (2014) and a sequence of important papers by Sun (2016), Malikov and Sun (2017), Sun and Malikov (2018) and Sun (2024), for example. More generally, semiparametric SAR models have attracted much interest, see e.g. Su and Jin (2010), Su (2012), Robinson and Thawornkaiwong (2012) and Zhang (2013), but their focus tends to be on inference on the parametric component.
@@
In this paper, we show how inference on varying coefficients in a range of SAR models can be conducted just like familiar parametric Wald tests for linear restrictions. The key idea is that if series approximations are used for the nonparametric varying coefficients then the series coefficients can be employed for very easy practical inference. This is motivated by the scientific aim of simplification of nonparametric/semiparametric inference. Indeed, a growing literature stresses such ideas, see e.g Ackerberg et al. (2012), Gupta (2018) and Korolev (2019), amongst others.
[…]
@@
Using this series expansion approach, we theoretically justify Wald-type test statistics for inference on varying coefficients. We cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence. The baseline model is similar to the one estimated by Sun et al. (2014), but also features a growing number of spatial lags and covariates. We then show how to allow for a degree of misspecification robustness by incorporating nonparametric spatial weights à la Pinkse et al. (2002). Our approach is also shown to work for models with varying coefficients on the spatial lags, in the spirit of Malikov and Sun (2017).
@@
We then apply our methods on two different studies where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This can be seen as a follow-up application to that of Li et al. (2002), who sugges that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e. output elasticities of capital and labor sum to unity) for most observed managerial expense levels. However, they do not have a test for CRS and do not model spatial dependence. Our analysis confirms their observations as we cannot reject the CRS hypothesis using their model and in much richer settings. This suggests that CRS technology may be a salient feature of this industry.
@@
Our second study is an illustration on the famous Boston housing data. We estimate a model similar to Sun et al. (2014), who used a model selection procedure to recommend that location should have a varying effect on house prices with respect to some other features of the property. We formally test for the varying location effects of these features using our statistics and indeed find them to be statistically significant.
@@
In a set of Monte Carlo simulation studies, we show that our tests have excellent finite-sample performance. In addition to tests that are based on the asymptotic normal approximation, we also propose a bootstrap version of our tests; the bootstrap tests perform particularly well for small sample sizes. We experiment with a range of SAR models and spatial dependence structures to demonstrate the full range of applicability and generality of our approach.
@@
The next section introduces our method in the context of a baseline higher-order SAR model. Section 3 extends the model to allow for spatial error dependence of Kelejian and Prucha (2007) type, while Section 4 allows for nonparametric spatial weights (Pinkse et al., 2002). In Section 5 we allow the coefficient on the spatial lag to also vary (Malikov and Sun, 2017) and show how our approach can be adapted to this setting. Section 6 presents our empirical applications to the Chinese nonmetal mineral industry and Boston house prices, while Section 7 contains the simulation study. Proofs are in the appendix, and a supplementary online appendix contains a number of additional simulation results.
2.   Baseline spatial autoregression
@@
where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance. Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow 0$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed. For all $i=1,\ldots,n$, $z_{in}$ is throughout the paper uncorrelated with $\epsilon_{in}$ while $x_{in}$ is allowed to be correlated with $\epsilon_{in}$.
Permitting $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modelling approach, with the model becoming richer as more observations become available to the practitioner. This approach has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference. Econometricians have also studied such models in least squares and GMM settings, see e.g. Andrews (1985), Koenker and Machado (1999) and Cattaneo et al. (2018), to name just three references. In a SAR setting, Gupta and Robinson (2015), Gupta (2018) and Gupta and Qu (2024) emphasize the advantages of this approach, especially when clustered data naturally imply asymptotic regimes in which $d_{\lambda}$ diverges with $n$. It is worth noting that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge.
We now mostly drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities. The paper takes the spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4. In all sections these weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004). This property is assumed to hold almost surely in Section 4, where the spatial weights are stochastic.
@@ 2.1   Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
@@
with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
@@
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$.
The null hypothesis to approximate $H_{0}^{true}$ is
@@
where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$, with $K_{1}$ being an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but having the same asymptotic order i.e. ${J_{1}}/{d_{\xi}}$ tends to a constant that is at least unity, as $n\rightarrow\infty$. Let $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\lambda}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix. . Then, the test statistic is
@@ 2.2   Asymptotic properties
We begin with restrictions on the moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$. Let $k_{ri}$ be the elements of the instrumental variable matrix, which corresponds to $K_{1},K_{2}$ and $K_{3}$ as defined in this section, 4 and 5, respectively, and $l_{ri}$ denote the elements of $L$.
@@
Assumption 1 is a general assumption that controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates for the approximation error because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 3 imposes moment conditions on regressors and instruments.
To ease notation, we introduce the following definition which captures boundedness and non-multicollinearity properties of random or fixed matrices of growing dimension.
@@
We take this opportunity to remind the reader that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
@@
Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem. This sequence of local alternatives features a $d_{\alpha}^{1/4}$ damping factor that accounts for the cost of our nonparametric approach, and has been found in similar problems by Hong and White (1995) and Gupta (2018), amongst others. We can now state the main theorem of the section.
Theorem 2.2.
@@
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{1}$ provides a consistent test.
[…]

### v2

@@ Abstract
We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of regression models allowing for cross-sectional or spatial dependence. We provide a general central limit theorem that covers a broad range of error spatial dependence structures, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial dependence parameters. Using our method, we first uncover evidence of constant returns to scale in the Chinese nonmetal mineral industry’s production function, and then show that Boston house prices respond nonlinearly to proximity to employment centers. A simulation study confirms that our tests perform very well in finite samples.
@@
Keywords: Spatial autoregression, varying coefficients, nonparametric inference
@@ 1.  Introduction
This paper develops a general framework for inference on varying coefficient regression models that is as easy to implement in practice as familiar Wald tests for finite-dimensional parameters. We also allow for spatial autoregressive (SAR) structure to permit interaction across the cross-section of economic agents. SAR models are a popular and parsimonious tool for researchers working in settings where economic agents interact via economic or social links. Such links can be geographic proximity, or proximity in some more general sense such as a social network. The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal, see e.g. Helmers and Patnam (2014) and Hsieh and van Kippersluis (2018) for examples of applications to peer effects. No surprise then that the baseline SAR model, due to Cliff and Ord (1981), has become much studied by econometricians. This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998), Lee (2004) and Kuersteiner and Prucha (2020).
@@
In keeping with the growing importance of SAR models and their ability to parsimoniously control for cross-sectional dependence in multiple regression designs, various strands of the literature have explored how to widen their scope. One such is the focus on varying coefficient SAR models, wherein model parameters are allowed to be unknown functions of some observed economic variable(s). This permits a flexible effect of both the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g. Fan and Zhang (1999, 2008)). In the SAR context, varying coefficient models have been studied by Sun et al. (2014) and a sequence of important papers by Sun (2016), Malikov and Sun (2017), Sun and Malikov (2018) and Sun (2024), for example. More generally, semiparametric SAR models have attracted much interest, see e.g. Su and Jin (2010), Su (2012), Robinson and Thawornkaiwong (2012) and Zhang (2013), but their focus tends to be on inference on the parametric component.
@@
In this paper, we show how inference on varying coefficients in a range of SAR models can be conducted just like familiar parametric Wald tests for linear restrictions. Standard varying coefficient regression without cross-sectional dependence is naturally covered as a special case. The key idea is that if series approximations are used for the nonparametric varying coefficients then the series coefficients can be employed for very easy practical inference. This is motivated by the scientific aim of simplification of nonparametric/semiparametric inference. Indeed, a growing literature stresses such ideas, see e.g Ackerberg et al. (2012), Gupta (2018) and Korolev (2019), amongst others. Naturally our methods are also applicable to the usual varying coefficient regression model with no cross-sectional dependence, see e.g. Ahmad et al. (2005), which is a simply a particular case of our theory.
To explain the intuition of our approach, suppose a sample of $n$ observations is available. With our method, the researcher can reduce the econometric problem of inference on an infinite-dimensional object to inference on a growing number of series coefficients, say $p$. In practice, this simply means writing down the usual Wald test statistic $\mathscr{W}$ for $p$ linear restrictions, observing that $p\rightarrow\infty$ as $n\rightarrow\infty$ if it is the length of a series approximation, and appealing to a theorem that establishes $(\mathscr{W}-p)/\sqrt{2p}\overset{d}{\rightarrow}N(0,1)$ as $n\rightarrow\infty$, under the null hypothesis. The idea stems from the fact that as $p\rightarrow\infty$, a $(\chi^{2}_{p}-p)/\sqrt{2p}$ random variable approaches a standard normal variate, see e.g. de Jong and Bierens (1994), Hong and White (1995), Gupta (2018) and Gupta and Seo (2023).
@@
Using this series expansion approach, we theoretically justify Wald-type test statistics for inference on varying coefficients. We first cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence. The baseline model is similar to the one estimated by Sun et al. (2014), but also features a growing number of spatial lags and covariates. We then show how to allow for a degree of misspecification robustness by incorporating nonparametric spatial weights à la Pinkse et al. (2002). Our approach is also shown to work for models with varying coefficients on the spatial lags, in the spirit of Malikov and Sun (2017).
@@
We then apply our methods in two different settings where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This can be seen as a follow-up study to Li et al. (2002), who suggest that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e. output elasticities of capital and labor sum to unity) for most observed managerial expense levels. Our analysis confirms their results as we cannot reject the CRS hypothesis using a more general version of their model that permits spatial dependence. This suggests that CRS technology may well be a salient feature of this industry.
@@
Our second study re-examines the Boston house price data of Harrison Jr and Rubinfeld (1978). We estimate a model similar to Sun et al. (2014), who used a model selection procedure to recommend that location should have a varying effect on house prices with respect to some other features of the property. We formally test for the varying location effects of these features using our statistics and indeed find them to be statistically significant.
@@
In Monte Carlo simulation studies, we experiment with a range of SAR models and spatial dependence structures to demonstrate the applicability and generality of our approach. Results show that our tests have excellent finite-sample performances.
@@
The rest of the paper is as follows: Section 2 introduces our method in a baseline regression model allowing for higher-order SAR structure. Section 3 extends the model to allow for spatial error dependence of the Kelejian and Prucha (2007) type. Section 4 allows for nonparametric spatial weights, as in Pinkse et al. (2002), and thereby permits some degree of weight matrix robustness. Section 5 shows how our approach can be adapted to the setting where the spatial lag coefficient varies, like in Malikov and Sun (2017). Section 6 applies our methods to the Chinese nonmetal mineral industry and Boston house prices. Section 7 presents the simulation study, and Section 8 concludes. Proofs are in the appendix, while additional empirical and simulation results are in the online appendix.
2.  Baseline higher-order spatial autoregression
@@
where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$-th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance. Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow\infty$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed. For all $i=1,\ldots,n$, $z_{in}$ is throughout the paper uncorrelated with $\epsilon_{in}$ while $x_{in}$ is allowed to be correlated with $\epsilon_{in}$.
It is worth noting that our results also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge. Allowing $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modeling approach, which has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference. Econometricians have also studied such models in least squares and GMM settings, see e.g. Andrews (1985), Koenker and Machado (1999) and Cattaneo et al. (2018), to name a few. In a SAR setting, Gupta and Robinson (2015) and Gupta (2018) emphasize the advantages of this approach, especially when clustered data imply asymptotic regimes where $d_{\lambda}$ diverges with $n$.
We now drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities. We take spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4. These weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004). This property is assumed to hold almost surely in Section 4, where the spatial weights are stochastic.
@@ 2.1 Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$ with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
@@
where $u_{i}=r_{i}+\epsilon_{i}$ and $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can then be written as
@@
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. The null to approximate $H_{0}^{true}$ is
@@
where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$ and $K_{1}$ is an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but the same asymptotic order, i.e. ${J_{1}}/{d_{\xi}}$ tends to a constant at least unity, as $n\rightarrow\infty$. Let $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\lambda}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix. Then, the test statistic is
@@ 2.2 Asymptotic properties
We begin with restrictions on moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$. Let $k_{ri}$ be elements of the instrumental variable matrix, corresponding to $K_{1},K_{2}$ and $K_{3}$ as defined in the following sections, and $l_{ri}$ denote the elements of $L$.
@@
Assumption 1 controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 3 imposes moment conditions on regressors and instruments.
To ease notation, we introduce the following definition that captures boundedness and non-multicollinearity of random or fixed matrices of growing dimension.
@@
We remind the reader that results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
@@
Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in detail in the proofs. This sequence of local alternatives features a $d_{\alpha}^{1/4}$ damping factor that accounts for the cost of our nonparametric approach, and has been found in similar problems by Hong and White (1995) and Gupta (2018), amongst others. We can now state the main theorem of the section.
Theorem $2.$
@@
as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{1}\overset{d}{\rightarrow}N(0,1)$. (ii) $\mathbb{W}_{1}$ provides a consistent test. (iii) Under the sequence of local alternatives $H_{\ell 1}$, $\mathbb{W}_{1}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.